# u-µP: The Unit-Scaled Maximal Update Parametrization

### v2

@@ 3.1 Not all training setups give µTransfer
We answer this Figure 2. The first training setup (a) is aligned with that used in Tensor Programs V (their Figure 4). There are several atypical aspects to their training setup, primarily the use of a constant LR schedule and a high number of epochs. This overfitting regime makes validation loss unusable, and transfer misleadingly good. When we remove these and shift to a standard Llama training setup (b), optimal HPs begin to drift with width. This confirms Lingle’s findings that standard µP is in fact a poor fit for modern LM training. We fix this (c) by the removal of parameters from LayerNorms/RMSNorms, as suggested by Lingle, and the introduction of independent weight decay for AdamW, as suggested by Wortsman et al. [18]  [^2] Lingle suggests independent weight decay is unstable, but we find it to be more so than Adam or standard AdamW. (see [19] for further analysis). With these changes adopted, we recover the strong transfer shown in Tensor Programs V’s experiments.
@@ 4.1 Combining µP with Unit Scaling
To maintain unit scale we apply $1/\sqrt{\mathrm{fan{\text{-}}in}}$ scaling in the backward pass (see Appendix H).
@@ 5.2 Quantifying hyperparameter interdependence
Our principled approach to HPs (Section 4.3) contains the requirement that their optimal values should depend minimally on the value of other HPs. We now investigate this empirically, conducting a 2D sweep over every pair of HPs for µP and u-µP, shown in Figures 9, LABEL: and 10 respectively.
@@ 5.4 Hyperparameter transfer
We demonstrate the transfer of LR across width in Figure 1 (b), of the other extended HPs across width in Figure 5, and of LR across training steps, batch size and depth in Figure 12. We find that:
@@
    The optimal LR is constant for all widths under u-µP, from 128 to 4096.
@@ A.1 Experimental Setup
A.2 Per-tensor learning rates
@@
A.3 Hyperparameter independence
In Section 5.2 we explore the question of HP independence under µP and u-µP. The following plots in Figures 9, LABEL: and 10 show the result of a 2D sweep over every pair of HPs under each scheme. All other HPs are held at 1 when not swept, except the $\eta$ which is held at $2^{-7.5}$ for µP and $2^{1.5}$ for u-µP, and $\hat{\eta}_{\mathrm{emb}}$ which is held at $2^{4}$ for µP.
These results show visual dependence between µP hyperparameters as a diagonal structure in the grids, such as $(\hat{\eta}_{\mathrm{emb}},\sigma_{\mathrm{init}})$ and $(\eta,\alpha_{\mathrm{attn}})$. We quantify this in the plot in Figure 4, where we use a measure of HP dependence termed transfer error. This is explained verbally in Section 5.2, and we provide an algorithmic description in Algorithm 1. We note that differences in transfer error between the two methods may also be influenced by the flatness of the optimum. The HP and loss values used for our transfer error calculations are those in Figures 9, LABEL: and 10.
@@ A.4 Hyperparameter search
A.5 Hyperparameter transfer experiments
@@ A.6 Numerical properties
A.7 Large-scale training details
@@ Appendix I From µP to u-µP
The default format used for training is the single-precision floating-point format, commonly known as FP32, with some hardware providers automatically casting it to the smaller TF32 compute mode for accelerated arithmetic. The 16-bit FP16 and BF16 formats were later introduced, and more recently the FP8 E5 & E4 formats [66, 67, 68]. The higher range of E5 has typically been used for gradients, while the higher precision of E4 has been seen as necessary for weights and activations. Our particular implementation of FP8 training is covered in LABEL:app:fp8_scheme. Other aspects of training such as the optimizer state and cross-device communication have also been put into FP8 [29], though not all tensors are amenable to being run in the lowest precision [40] without degradation. The use of multiple formats is known as mixed precision [69]. A comparison of these formats is given in Table 11.

### v3

@@ 3.1 Not all training setups give µTransfer
We answer this in Figure 2. The first training setup (a) is aligned with that used in Tensor Programs V (their Figure 4). There are several atypical aspects to their training setup, primarily the use of a constant LR schedule and a high number of epochs; we outline the precise differences between setup (a) and (b) in Table 6. This overfitting regime makes validation loss unusable, and transfer misleadingly good. When we remove these and shift to a standard Llama training setup (b), optimal HPs begin to drift with width (see Figure 9 for an ablation). This confirms Lingle’s findings that standard µP is in fact a poor fit for modern LM training. We fix this (c) by the removal of parameters from LayerNorms/RMSNorms, as suggested by Lingle, and the introduction of independent weight decay for AdamW, as suggested by Wortsman et al. [18]  [^2] Lingle suggests independent weight decay is unstable, but we find it to be more so than Adam or standard AdamW. (see [19] for further analysis). With these changes adopted, we recover the strong transfer shown in Tensor Programs V’s experiments.
@@ 4.1 Combining µP with Unit Scaling
To maintain unit scale we apply $1/\sqrt{\mathrm{fan{\text{-}}out}}$ scaling in the backward pass (see Appendix H).
@@ 5.2 Quantifying hyperparameter interdependence
Our principled approach to HPs (Section 4.3) contains the requirement that their optimal values should depend minimally on the value of other HPs. We now investigate this empirically, conducting a 2D sweep over every pair of HPs for µP and u-µP, shown in Figures 14, LABEL: and 15 respectively.
@@ 5.4 Hyperparameter transfer
We train many models and plot transfer of LR across width (Figure 1 (b)), steps, batch size and depth (Figure 5), and transfer of other HPs across width (Figure 17). Note that u-µP (building on µP) is designed to give transfer over width [^4] As we use depth-µP this could be said about depth as well, but as [14] show that transformers don’t attain depth-transfer under depth-µP we do not expect strong transfer across depth.; the other axes we report for practical purposes. We find that:
@@
    The optimal LR is constant across width under u-µP. There is a small drift for training steps and batch size, and a larger one with depth. Hence we recommend proxy models which primarily differ in width, moderately in steps and batch size, and least in depth.
@@ A.1 Experimental Setup
A.2 Further analysis of µTransfer failure modes
In Table 6 we provide exact details of the experimental differences between setups (a) and (b) from Figure 2, for readers wishing to understand and reproduce this result. We also provide a step-by-step ablation of the various changes made between these setups in Figure 9.
For setup (c), which shows how our two combined stability fixes mitigate the problem of poor transfer, both changes are evaluated independently in Figure 8, which shows that the dominant effect is a narrowing of the learning basin due to non-parametric RMSNorms, leading to better learning rate transfer.
Figure 8: The effect of the individual transfer stability fixes from Figure 2. (a) In this setting switching from non-independent to independent weight decay has only a minor effect, though [18] Figure 6 suggests it may be highly valuable in other settings. (b) Non-parametric norms give a narrower learning rate basin, leading to better transfer. (c) The combination of these, for comparison, matching Figure 2 (c).
Table 6: Comparison of Tensor Programs V’s standard settings (as best we can tell) and our Standard Llama setup, corresponding to (a) and (b) in Figure 2.
Feature Tensor Programs V Standard Llama
Dataset wikitext-2 wikitext-103
Vocab Size 33278 32000
Nsteps 10000 8192
Batch Size 20 64
Optimizer adam adamw
LR Schedule constant cosine
Weight Decay 0 0.00012
Positional Encoding absolute rotary
Norm layer_norm rms_norm
Dropout 0.2 0
NLayers 2 4
Use Gated FFN False True
Activation FN relu swish
FFN Ratio 4 2.75
Final Norm False True
Base Depth 1 4
Zero Init Readout True False
Figure 9: An ablation of the more standard Llama training settings against the Tensor Programs V settings from Figure 2. This shows that the flat basins with poor transfer are not due to a single change, but the combination of a larger dataset (training $<\!1$ epoch) and the stronger Llama model are largely responsible. Note that ‘Llama model’ here indicates a group of changes: rms norm, rotary embeddings & swiglu FFN.
A.3 Validating our experimental setup
In this section we run a series of ablations to validate decisions relating to our experimental setup given above. In particular, we examine the effect of using repeated data, the effect of using a shorter warmup duration, and the effect of different final learning rates at the end of decay.
A.3.1 Repeated data
As outlined in Table 5, our standard training setup uses 0.97 epochs of the WikiText-103 dataset (50x larger than the WikiText-2 dataset used in Tensor Programs V). However on our batch size and training steps scaling experiments in Figure 5 we train on up to 4×4$\times$4 × the amount of data than in our standard setup, and hence use up to $4$ epochs.
Though this is still a small level of repeated data, this moves our training slightly into the over-fitting regime. Based on this change, we here investigate the hypothesis that this regime has better or worse transfer of the optimal LR than the non-overfitting regime, and hence our results could be misleading. To do so, we repeated these experiments with the same number of tokens, but using the much larger SlimPajama dataset [23] where we use $<1$ epoch.
The results for this experiment are seen in Figure 10. The shape of curves is very similar across the two datasets, for both batch size and training steps (albeit with a higher loss, due to the more varied nature of SlimPajama). From this we conclude that the effect of repeated data from our use of WikiText-103 is not significant.
Figure 10: A repeat of the batch size and training steps experiments in Figure 5, but using the larger SlimPajama dataset where no data is repeated. In both settings our validation loss basins take the same shape, indicating that our analysis using the WikiText-103 dataset holds.
A.3.2 Warmup duration
For our experimental setup (Table 5) we use a longer duration of warmup than in our large-scale setup (Table 7). We do so out of caution, as we use fewer tokens-per-batch for the smaller-scale experiments and so may require longer warmup. However, doing so also creates the risk of spending too large a proportion of training doing warmup, which could affect transfer.
To investigate this effect, we run two experiments. Firstly, we re-run the experiment for LR transfer over training steps, shown in Figure 5, on a quarter of the warmup steps. This is shown in Figure 11 (left). The main effect appears to be higher loss for larger learning rates, but the optima are largely unchanged. The only exception is the 4096-step run, where the optimum shifts left and the loss improves slightly. This appears to now align the optimum better with the other training durations, but leads to narrower basins as a result, suggesting a trade-off for this particular experiment.
However, all our other experimental runs use the 8192-step configuration, which has a consistent optimum regardless of warmup duration here. To investigate the effect of reduced warmup on width transfer at this particular step-count, we re-run our experiment in Figure 1 (b) under the shorter warmup duration, shown in Figure 11 (right). The only significant impact of this change is to narrow the basins, inducing no significant change in the optimal LR. As such, we conclude that using 2000 steps of warmup in our experimental setup is a reasonable choice, and both give the same width transfer.
Figure 11: (Left) Learning rate transfer across training steps under different numbers of warmup steps. (Right) Learning rate transfer across width under different numbers of warmup steps. In this setting (training steps = 8192) the optimal LR is consistent, meaning either warmup regime can be used, though the longer gives wider basins.
A.3.3 Learning rate decay target
In all our experiments we use a cosine decay of our learning rate down to 10% of the maximum. This follows the standard approach taken by most LLM training projects [50, 24, 51, 10, 52]. However, recent research has indicated that this may not be the optimal decay target, with implications for LR transfer. [53] show that the choice of target percentage can alter the shape of transfer curves and potentially shift the optimum value (Figure 21, right). They also suggest that using a fixed target value may work better than a percentage (Figure 22, right), which could be swept separately. [54] separately suggest that linear decay to zero is the most effective scheme.
Though using the optimal decay scheme is not necessarily essential to the validity of our method, any implications of different schemes on transfer properties should be investigated. To do so, we run two experiments. The first sweeps the learning rate for our standard model at various percentages and fixed values of cosine decay target, including zero, in Figure 12 (left). Lower decay targets perform better here, including zero, suggesting that this simple rule may be ideal.
We then re-run our width transfer experiment from Figure 1 (b) but with our LR decaying to 0, and plot the result in Figure 12 (right). This leads to slightly better results for large learning rates, though for large models this difference diminishes. Fortunately the effect this decay target has on the shape of curves (and hence optimal LR transfer) is minimal, indicating that our conclusions are not effected significantly by the choice of decay target.
Figure 12: (Left) A learning rate sweep over LR targets of different types (percentage, fixed and zero) on our standard model. (Right) Using the zero and 10% learning rate targets, LR transfer over width.
A.4 Per-tensor learning rates
@@
A.5 Hyperparameter independence
In Section 5.2 we explore the question of HP independence under µP and u-µP. The following plots in Figures 14, LABEL: and 15 show the result of a 2D sweep over every pair of HPs under each scheme. All other HPs are held at 1 when not swept, except the $\eta$ which is held at $2^{-7.5}$ for µP and $2^{1.5}$ for u-µP, and $\hat{\eta}_{\mathrm{emb}}$ which is held at $2^{4}$ for µP.
These results show visual dependence between µP hyperparameters as a diagonal structure in the grids, such as $(\hat{\eta}_{\mathrm{emb}},\sigma_{\mathrm{init}})$ and $(\eta,\alpha_{\mathrm{attn}})$. We quantify this in the plot in Figure 4, where we use a measure of HP dependence termed transfer error. This is explained verbally in Section 5.2, and we provide an algorithmic description in Algorithm 1. We note that differences in transfer error between the two methods may also be influenced by the flatness of the optimum. The HP and loss values used for our transfer error calculations are those in Figures 14, LABEL: and 15.
@@ A.6 Hyperparameter search
A.7 Hyperparameter transfer experiments
@@ A.8 Numerical properties
A.9 Large-scale training details
@@ Appendix I From µP to u-µP
The default format used for training is the single-precision floating-point format, commonly known as FP32, with some hardware providers automatically casting it to the smaller TF32 compute mode for accelerated arithmetic. The 16-bit FP16 and BF16 formats were later introduced, and more recently the FP8 E5 & E4 formats [71, 72, 73]. The higher range of E5 has typically been used for gradients, while the higher precision of E4 has been seen as necessary for weights and activations. Our particular implementation of FP8 training is covered in Section 4.2. Other aspects of training such as the optimizer state and cross-device communication have also been put into FP8 [29], though not all tensors are amenable to being run in the lowest precision [40] without degradation. The use of multiple formats is known as mixed precision [74]. A comparison of these formats is given in Table 12.
@@
Given that the end-goal of leveraging u-mup’s low-precision properties is to speed up training and reduce memory usage, it’s reasonable to ask why we don’t investigate this experimentally. The answer relates to the relative immaturity of the FP8 training software stack - a lack of open, efficient FP8 kernels for compute and communication mean significant additional engineering effort is required to attain expected speedups over the full model.
Here we show that u-µP’s static scaling factors add no overhead to matmuls in FP8, and hence ought to be able to reach close to the maximal FP8 throughput attainable for the full model.
@@
The above analysis focuses on throughput; significant memory savings are also possible through the use of FP8, though how this affects the total memory footprint depends on various additional variables and the overall distributed training setup. The following factors are play a significant role: typically the main memory bottlenecks are the optimizer states, which are kept in full precision. This footprint can be reduced by applying ZeRO sharding [55], though for significant gains the number of data parallel processes needs to be sufficiently large and ZeRO stage 2 or 3 are required. In these settings the memory footprint of activations and gradients becomes significant, and quantizing these to lower precision promises further memory savings, though may be non-trivial [29].